OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (solomonzelman):

Divide both sides by -3.

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

|x+7|=4 this implies x+7=4, x+7=-4 this implies x=-3,-11

OpenStudy (anonymous):

thank you i understand now.